OpenStudy (anonymous):

x^2 - 2x -24 ______________ x-6

geerky42 (geerky42):

HINT: x²-2x-24 = (x-6)(x+4)

OpenStudy (anonymous):

\[\frac{ x ^{2}-2x-24 }{ x-6 }\]

OpenStudy (anonymous):

divide